# Node classification in networks via simplicial interactions

### v2

@@ III-A Simplicial objective function
Figure 1: Overall optimization process. First, we classify simplices in a given network based on their sizes (Breaking Down). Then, we multiply each size category by a distinct multinomial coefficient, sum these weighted values, and use this combined information to train a model (Training). Finally, we evaluate the model’s classification performance (Classification).
@@ III-B SBTM: a new graph generation approach
However, SBM has notable limitations: it lacks robust local clustering, failing to capture dense substructures like triangles essential for representing complex network dynamics. SBMs also assume edge independence, which is unrealistic as actual networks exhibit interdependencies among edges and a mix of hierarchical community structures [27]. Consequently, SBMs may not accurately reflect the diverse characteristics and higher-order interactions commonly observed in real-world networks. To address this, in this paper we present the Stochastic Block Tensor Model (SBTM), which generalizes SBM by assigning probability tensors to generate $k$-cliques for each $k\geq 2$, facilitating the development of higher-order simplices in the network. This results in the formation of higher-order simplices in the synthetic network, yielding networks that closely mimic real-world networks.
@@ V-C Integration of the objective function with GNN
In this experiment, we integrate GAT [22] with the proposed objective function and evaluate the performance gains using the Cora, Citeseer, and Pubmed datasets. GAT uses an attention mechanism to learn node embeddings. The node features are created using the bag-of-words representation of documents, with the dimensions of the node features for Cora, Citeseer, and Pubmed being 1433, 3703, and 500, respectively. These features are used to learn embedding vectors using a multi-head attention structure, and a linear map is employed to generate outputs corresponding to the number of clusters. In our method, the total training and validation data used in [22]—640 for Cora, 620 for Citeseer, and 560 for Pubmed—are treated as the overall prior information. In contrast to [22], which evaluated performance using 1,000 test nodes for each dataset, we assess performance on all remaining nodes not designated for training or validation. According to the results of integration experiments, the accuracies using training and validation nodes in the GAT study as prior information (label-aware nodes) are 84.9%, 72.9%, and 80.1% for Cora, Citeseer, Pubmed dataset respectively (Table II).
@@ V-D Summary and discussion
Sixth, according to experimental results with benchmark data, using the training and validation data employed in GAT research [22] as prior information improves accuracy by 1.9%, 0.4%, and 1.1% for Cora, Citeseer, and Pubmed, respectively, compared to the result in [22]. The objective function used in this study is intended to promote that all nodes forming higher-order simplices exhibit similar node probability distributions. As a result, if the network has a large number of inter-simplices, where nodes producing higher-order simplices belong to distinct labels or clusters, achieving good performance with the proposed objective function becomes challenging. The proportions of inter-cluster simplices among all higher-order simplices beyond 2-simplices are 16.4%, 18.8%, and 22.6% (therefore, intra-cluster simplices ratio would be 83.6%, 81.2%, 77.4%) for Cora, Citeseer, and Pubmed, respectively. This observation explains why a higher accuracy improvement is observed with Cora compared to Citeseer and Pubmed.

### v3

@@ III-A Simplicial objective function
Figure 1: Overall optimization process. First, we classify $k$-simplices (for $k=1,2,…$) in a given network based on their sizes (Breaking Down). Then, we multiply each size category by a distinct multinomial coefficient, sum these weighted values, and use this combined information to train a model (Training). Finally, we evaluate the model’s classification performance (Classification).
@@ III-B SBTM: a new graph generation approach
However, SBM has notable limitations: it lacks strong local clustering, failing to capture dense substructures like triangles that are essential for representing complex network dynamics. Additionally, SBMs assume edge independence, which is unrealistic since real-world networks often exhibit interdependencies among edges along with a combination of hierarchical community structures [28]. When comparing networks of equal density—networks with an identical number of nodes and edges—SBM-based models generally yield far fewer high-dimensional polyhedrons (simplifies or cliques) than what is commonly observed in real-world networks. This shortcoming arises from SBM’s foundation as an edge-generation model, which limits its ability to form interconnected groups beyond pairs. For example, in collaboration networks where multiple individuals contribute to the same project or publication, groups of three or more people often establish simultaneous collaborative connections. Such higher-order connections are challenging for traditional SBMs to capture accurately. Consequently, edge-generation models like SBM struggle to recreate the complex, interdependent structures present in real-world networks. To address this, in this paper we present the Stochastic Block Tensor Model (SBTM), a natural extension of SBM by assigning probability tensors (rather than a single probability matrix) to generate $k$-cliques for each $k\geq 2$, facilitating the development of higher-order simplices in the network. This leads to the emergence of higher-order simplices in the synthetic network, resulting in networks that closely resemble real-world networks.
@@ V-C Integration of the objective function with GNN
In this experiment, we integrate GNNs with the proposed objective function and evaluate the performance gains using the Cora, Citeseer, and Pubmed datasets. GAT [22] uses an attention mechanism to learn node embeddings. The node features are created using the bag-of-words representation of documents, with the dimensions of the node features for Cora, Citeseer, and Pubmed being 1433, 3703, and 500, respectively. These features are used to learn embedding vectors using a multi-head attention structure, and a linear map is employed to generate outputs corresponding to the number of clusters. In our method, the total training and validation data used in [22]—640 for Cora, 620 for Citeseer, and 560 for Pubmed—are treated as the overall prior information. In contrast to [22], which evaluated performance using 1,000 test nodes for each dataset, we assess performance on all remaining nodes not designated for training or validation. The treatment is similarly applied to Planetoid [41] (transductive experiment), GCN [20] (with 64 hidden units), and SGC [50]. Table II summarizes the results of the integration experiments.
@@ V-D Summary and discussion
Sixth, according to experimental results with benchmark data, using the training and validation data employed in Planetoid, GCN, GAT, and SGC as prior information improves mean accuracy by 2.2%, 0.7%, and 0.3% for Cora, Citeseer, and Pubmed, respectively. The objective function used in this study is intended to promote that all nodes forming higher-order simplices exhibit similar node probability distributions. As a result, if the network has a large number of inter-simplices, where nodes producing higher-order simplices belong to distinct labels or clusters, achieving good performance with the proposed objective function becomes challenging. The proportions of inter-cluster simplices among all higher-order simplices beyond 2-simplices are 16.4%, 18.8%, and 22.6% (therefore, intra-cluster simplices ratio would be 83.6%, 81.2%, 77.4%) for Cora, Citeseer, and Pubmed, respectively. This observation explains why a higher accuracy improvement is observed with Cora compared to Citeseer and Pubmed.
Seventh, in widely used real datasets such as Cora, Citeseer, and Pubmed, the number of labels $l$ is modest, allowing the proposed algorithm’s time complexity to remain manageable, and our method can be effectively applied without significant computational concerns. However, for datasets with a large number of labels, such as the Amazon Product Co-purchasing Network [59], which contains hundreds or even thousands of labels, the computational cost may be too high. Because our algorithm’s complexity increases exponentially with $l$, it may be impractical to utilize it on datasets with a large number of labels. This constraint highlights the need for algorithmic improvements to support applications with multiple labels.